OpenStudy (anonymous):

Find the midpoint of the segment with the given endpoint. D(1, 2) and E(-3, 6). option are- A. (-7, 10) B. (-2, 2) C. (-1, 4) D. (4, 64)

OpenStudy (anonymous):

To find the mid-point we need to find the difference between the x-values of the 2 points and then half it, and do the same for the y values. Are you able to do this now?

OpenStudy (raden):

midpoint is avarage points coordinat

OpenStudy (anonymous):

Technically, if we have 2 points: \[(a, b)\] and \[(c, d)\] Then the midpoint of the two will be: \[(\frac{a+c}{2}, \frac{b+d}{2})\]

OpenStudy (raden):

xm = (x1+x2)/2 ym = (y1+y2)/2